OpenStudy (anonymous):

1/8%, express percent as a decimal.

OpenStudy (anonymous):

0.125% ? i think

OpenStudy (anonymous):

1/8%=0.125% 0.00125 is the decimal form.

OpenStudy (anonymous):

1/8=.125 and .125*100= 12.5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

amit7808, it is not 12.5 because that won't give us 1/8%.

OpenStudy (anonymous):

hope u understood wel to award medel :)

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

Hi Pat18 it will be same http://www.mathsisfun.com/decimal-fraction-percentage.html

OpenStudy (anonymous):

to support my answer

OpenStudy (anonymous):

nope. it won't be the same.

OpenStudy (anonymous):

actually it is not, the given is not 12.5%, it is 0.125%, hope you observe it.

OpenStudy (anonymous):

i don't need the website amit7808.

OpenStudy (anonymous):

amit7808 this is the process, if we are going to change percent to decimal, all we have to do is to move the decimal point two times to the left.

OpenStudy (anonymous):

yeah the answer is 0.125% if you derive it into decimal form you will get .00125 that is just how simple it is. :) peace out guys!

OpenStudy (anonymous):

I misunderstood the problem .. thnx Pat and bio :)

OpenStudy (anonymous):

thank you for accepting the mistake. peace:)

OpenStudy (anonymous):

okiez...amit7808 ^_^